OpenStudy (anonymous):

an aircraft weighing 6400 pounds lands at a speed of 10ft/sec and stops in 10 secs. what force was generated by the brakes (assuming gravity as 32ft/sec).

OpenStudy (anonymous):

without friction? \[F_{e} = \sqrt{[(ma)^2 + (mg)^2] }\] \[a = \Delta v /\Delta t = 10 / 10 [ft/\sec/\sec] = 1 ft / \sec^2 \] \[F_{e} = \sqrt{(6400*1)^2 + (6400*32)^2} = 204899.97 [pounds*ft/\sec^2]\]